OpenStudy (anonymous):

To get from your house to your grandparent’s house you must avoid walking through a pond. To avoid the pond, you walk 0.5 miles south and 1.2 miles east. How many miles would be saved if it were possible to walk through the pond? A. 0.4 miles B. 1.69 miles C. 3.0 miles D. 6.4 miles

OpenStudy (anonymous):

It will be a right angled triangle ..

OpenStudy (anonymous):

With two sides 1.2 & .5 . Square each . then Add then take sqrt

OpenStudy (anonymous):

2.25+0.25=6.25 @Anu2401

OpenStudy (anonymous):

Yes take sqrt of 6.25. This will be 2.5 Answer. In such a situation always draw a rough diagram. see attachment :)

OpenStudy (anonymous):

the answer is C @Anu2401

OpenStudy (anonymous):

Oh its (1.2)^2 +(.5)^2 = 1.69 Then Sqrt of 1.69 is 1.3 Hence length of is 1.3

OpenStudy (anonymous):

Then saving will be 1.2+.5-1.3=0.4 is the correct one

OpenStudy (anonymous):

See

OpenStudy (anonymous):

C is not the correct answer @Ambbiiee its A

OpenStudy (anonymous):

Is this OK r u getting it @ @Ambbiiee

OpenStudy (anonymous):

i am confused around the part where you started √69

OpenStudy (anonymous):

@Anu2401

OpenStudy (anonymous):

its sqrt(1.69)

OpenStudy (anonymous):

Now is it OK ??

OpenStudy (anonymous):

yeah thanks

OpenStudy (anonymous):

Welcome :)